13) Suppose rn =2, πT=2, ay =1 and aπ =1. Suppose an adverse supply shock in year 0 causes inflation to jump to 4%. Assume the usual one-year lag on output and another year lag for inflation. If the central bank follows the Taylor Rule then we would expect r0 = ___ , r1 = _____ and r2 = _____.

Taylor Equation r = rn + ayỸ+ aπ(π πT)

a) 4;3;2

b) 3;4;2

c) 4;2;2

d) 3;3;2
Answer #1

Hi! Welcome to Chegg!

Taylor Equation r = rn + ayỸ+ aπ(π πT)

Due to one year lag in output, initial output stays constant.

So, r0 = 2 + 1(0) + 1(4-2) = 4

After 1 year, effect on output comes but inflation stays constant at 4% (as given one year lag).

So, r1 = 2 + 1(-1) + 1(4-2) = 3

After 2 years, output takes previous year value but inflation changes now.

So, r2 = 2 + 1(-1) + 1(3-2) = 2

Ans (a): 4;3;2